# Partially-Supervised Neural Network Model For Quadratic Multiparametric Programming

### v1

@@ 5 Numerical Results
Optimal Power Flow (OPF) is a central problem in electrical power system management [28]. The problem is to find the optimal dispatch per generator to minimize an objective function (e.g., cost of operation) while satisfying energy demands and grid system constraints at a given point in time. The formulation of the AC-OPF (alternating current) problem is nonlinear and nonconvex, so challenges in finding optimal and feasible solutions led researchers to use the simplified DC-OPF (direct current) formulation which has linearized power flow constraints under certain assumptions. The DC-OPF problem can be defined as:
@@ 5.1 Test Results
Our method was evaluated using IEEE-6, IEEE-30, IEEE-57 bus test systems implemented in MATPOWER package LABEL:zimmerman2010matpower. We used our PSNN model to predict optimal solutions to the DC-OPF problem, and compared them to solutions predicted by a NN model trained using classical methods (i.e., with paired input-output data representing particular values of system parameters and corresponding optimal solutions), and also to solutions generated analytically by the Gurobi solver (v10.0.3), which provides a baseline experimental control for comparing performance of the NN models. The PSNN models were trained using only 1000 data points, and validated and tested on datasets of 1000 and 4000 observations, respectively. To generate training and validation data, demand of all buses was set to 0.01, then one bus was chosen and its load incrementally increased from 0 to the maximum allowed by the system, and the procedure was repeated for one other bus. For the classically trained NN model, we generated training, validation and test sets of 5000, 1000 and 4000 observations, respectively, by setting the demand to the default load values in the Pandapower package [29], multiplying with a constant sampled from Uniform$(0.6,1.4)$ and solving the problem using Gurobi solver for each generated demand value [8, 30]. All models were tested using two types of test datasets, generated to reflect realistic and extreme ranges of demand. Realistic demand data was generated by multiplying a base demand by a constant sampled from Uniform$(0.6,1.4)$. Extreme demand data was generated by fixing all load bus demands to 0.01, replacing one bus with a number sampled from $Uniform(0,\max)$ and repeating this step for all buses. All the experiments and training were done on a Mac Mini M2 (2023) on Python 3.9.18, using PyTorch 2.0.1.
@@
KKT 1 $(\textbf{P}_{g})$ KKT 1 $(\boldsymbol{\delta})$ KKT 2 KKT 2 ()($\leq$)( ≤ ) KKT 3 KKT 4
@@
KKT1-$\textbf{P}_{g}$ KKT1-$\boldsymbol{\delta}$ KKT2 KKT2 ()($\leq$)( ≤ ) KKT3 KKT4
@@
KKT1-$\textbf{P}_{g}$ KKT1-$\boldsymbol{\delta}$ KKT2 KKT2 ()($\leq$)( ≤ ) KKT3 KKT4
@@ 5.2 Uncertainty
where $\textbf{P}_{ren}\sim$ Exponential$(\lambda=1.25)$. In the simulations, the problem was solved repeatedly for 500 random cases of $\textbf{P}_{ren}$ for each hour. A 1.5 unit generator limit was enforced by truncating the generated values above 1.5.
@@
For this analysis, default load values from the Pandapower package [29] were taken as base demand. To reflect seasonality of hourly changing demand, we used historical demand data for Ontario, Canada  [^2] https://www.ieso.ca/en/power-data/data-directory as a scaler. Specifically, an arbitrary day (May 11, 2024) was selected and the daily demand was divided by the maximum value throughout the day to act as scalers. Each bus value was multiplied by the constant for every hour to obtain 24 demand inputs. To calculate solutions for 1200 different inputs, the PSNN model ran for 0.03 seconds.
@@ A.1 Neural Networks as Piecewise Linear Functions
$\begin{cases}0&\text{ for }\textbf{x}\in\mathcal{R}_{0}\\ \sum_{j\in\mathcal{B}_{1}}\textbf{W}^{1}_{j}\textbf{W}^{0}_{j}x_{i}+\textbf{W}^{1}_{j}b^{0}_{j}&\text{ for }\textbf{x}\in\mathcal{R}_{1}\\ \ \vdots&\vdots\\ \sum_{j\in\mathcal{B}_{K-1}}\textbf{W}^{1}_{j}\textbf{W}^{0}_{j}x_{i}+\textbf{W}^{1}_{j}b^{0}_{j}&\text{ for }\textbf{x}\in\mathcal{R}_{K-1}\\ \sum_{j\in\mathcal{B}_{K}}\textbf{W}^{1}_{j}\textbf{W}^{0}_{j}x_{i}+\textbf{W}^{1}_{j}b^{0}_{j}&\text{ for }\textbf{x}\in\mathcal{R}_{K}\\ \end{cases}.$ (21)
@@
$f(x_{i},\theta)=\begin{cases}\textbf{N}_{0}&\text{ for }\textbf{x}\in\mathcal{R}_{0}\\ \textbf{M}_{1}\textbf{x}+\textbf{N}_{1}&\text{ for }\textbf{x}\in\mathcal{R}_{1}\\ \ \vdots&\vdots\\ \textbf{M}_{K-1}\textbf{x}+\textbf{N}_{K-1}&\text{ for }\textbf{x}\in\mathcal{R}_{K-1}\\ \textbf{M}_{K}\textbf{x}+\textbf{N}_{K}&\text{ for }\textbf{x}\in\mathcal{R}_{K}\\ \end{cases}.$ (22)
@@
Figure 12 illustrates how each activated neuron changes the slope of the piecewise linear function represented by a NN. The points donated as $Act.\mathcal{B}_{1}$ and $Act.\mathcal{B}_{2}$ show where different sets of binding constraints are activated and NN diagrams on the left show activated neurons at respective points. As illustrated, for $x_{3}<2$, none of the hidden neurons have a positive value as the $\textbf{W}_{j}^{0}x+b_{j}<0$ for all $j$. In the interval $42\leq x_{3}<$ 4, the set of active neurons is $\mathcal{B}_{1}=\{1\}$ as only $\textbf{W}_{1}^{0}x+b_{1}>0$ and the output increases with a certain slope as $x_{3}$ increases. After $x_{3}\geq 4$, another slope is also activated and the set of active slopes are $\mathcal{B}_{2}=\{1,5\}$, which updates the change of the output value with respect to $x_{3}$.
@@ A.4 Further Details on Training of PSNN
While PSNN was tested on these datasets, the training set is constructed using the critical regions $[\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ detected via the discovery step. To discover sets of binding constraints, all parameters were set to 0.01 and an arbitrary $\theta_{i}^{\prime}$ was incrementally increased from 0 to the maximum number the system allows. The discovery was carried out on one dimension only, so this approach would not detect if other sets of binding constraints could be found by repeating this step with other $\theta_{i}$.
To construct the training set for PSNN, different values of $\theta_{i}^{\prime}$ were sampled from each $[\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ and solved by expanding the coefficient matrix, $\textbf{J}_{\mathcal{B_{i}}}$ with the corresponding binding constraints, running the calculations in eq. 13 and checking if the solutions satisfy the KKT conditions. This step was repeated using one other $\theta_{j}$ and fixing all other parameters, assuming the same $[\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ would apply to this range. If the calculated solution does not satisfy KKT, another number was sampled from the same interval.

### v2

@@ 5 Numerical Results
Optimal Power Flow (OPF) is a central problem in electrical power system management (carpentier1979optimal, ). The problem is to find the optimal dispatch per generator to minimize an objective function (e.g., cost of operation) while satisfying energy demands and grid system constraints at a given point in time. The formulation of the AC-OPF (alternating current) problem is nonlinear and nonconvex, so challenges in finding optimal and feasible solutions led researchers to use the simplified DC-OPF (direct current) formulation which has linearized power flow constraints under certain assumptions. The DC-OPF problem can be defined as:
@@ 5.1 Test Results
Our method was evaluated using IEEE-6, IEEE-30, IEEE-57 bus test systems implemented in MATPOWER package zimmerman2010matpower . We used our PSNN model to predict optimal solutions to the DC-OPF problem, and compared them to solutions predicted by a NN model trained using classical methods (i.e., with paired input-output data representing particular values of system parameters and corresponding optimal solutions), and also to solutions generated analytically by the Gurobi solver (v10.0.3), which provides a baseline experimental control for comparing performance of the NN models. The PSNN models were trained using only 1000 data points, and validated and tested on datasets of 1000 and 4000 observations, respectively. To generate training and validation data, demand of all buses was set to 0.01, then one bus was chosen and its load incrementally increased from 0 to the maximum allowed by the system, and the procedure was repeated for one other bus. For the classically trained NN model, we generated training, validation and test sets of 5000, 1000 and 4000 observations, respectively, by setting the demand to the default load values in the Pandapower package thurner2018pandapower , multiplying with a constant sampled from Uniform(0.6,1.4)(0.6,1.4) and solving the problem using Gurobi solver for each generated demand value zamzam2020learning ; joswig2022opf . All models were tested using two types of test datasets, generated to reflect realistic and extreme ranges of demand. Realistic demand data was generated by multiplying a base demand by a constant sampled from Uniform(0.6,1.4)(0.6,1.4). Extreme demand data was generated by fixing all load bus demands to 0.01, replacing one bus with a number sampled from $Uniform(0,\max)$ and repeating this step for all buses. All the experiments and training were done on a Mac Mini M2 (2023) on Python 3.9.18, using PyTorch 2.0.1.
@@
KKT 1 $(\textbf{P}_{g})$ KKT 1 $(\boldsymbol{\delta})$ KKT 2 KKT 2 $(\leq)$ KKT 3 KKT 4
@@
$\textbf{P}_{g} KKT1$-$\boldsymbol{\delta}$ KKT2 KKT2 $(\leq)$ KKT3 KKT4
@@
$\textbf{P}_{g} KKT1$-$\boldsymbol{\delta}$ KKT2 KKT2 $(\leq)$ KKT3 KKT4
@@ 5.2 Uncertainty
[…]
@@
For this analysis, default load values from the Pandapower package thurner2018pandapower were taken as base demand. To reflect seasonality of hourly changing demand, we used historical demand data for Ontario, Canada  [^2] https://www.ieso.ca/en/power-data/data-directory as a scaler. Specifically, an arbitrary day (May 11, 2024) was selected and the daily demand was divided by the maximum value throughout the day to act as scalers. Each bus value was multiplied by the constant for every hour to obtain 24 demand inputs. To calculate solutions for 1200 different inputs, the PSNN model ran for 0.03 seconds.
@@ A.1 Neural Networks as Piecewise Linear Functions
[…]
@@
[…]
@@
Figure 12 illustrates how each activated neuron changes the slope of the piecewise linear function represented by a NN. The points donated as $Act.\mathcal{B}_{1}$ and $Act.\mathcal{B}_{2}$ show where different sets of binding constraints are activated and NN diagrams on the left show activated neurons at respective points. As illustrated, for $x_{3}<2$, none of the hidden neurons have a positive value as the $\textbf{W}_{j}^{0}x+b_{j}<0$ for all $j$. In the interval $2\leq x_{3}<4$, the set of active neurons is $\mathcal{B}_{1}=\{1\}$ as only $\textbf{W}_{1}^{0}x+b_{1}>0$ and the output increases with a certain slope as $x_{3}$ increases. After $x_{3}\geq 4$, another slope is also activated and the set of active slopes are $\mathcal{B}_{2}=\{1,5\}$, which updates the change of the output value with respect to $x_{3}$.
@@ A.4 Further Details on Training of PSNN
While PSNN was tested on these datasets, the training set is constructed using the critical regions $\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ detected via the discovery step. To discover sets of binding constraints, all parameters were set to 0.01 and an arbitrary $\theta_{i}^{\prime}$ was incrementally increased from 0 to the maximum number the system allows. The discovery was carried out on one dimension only, so this approach would not detect if other sets of binding constraints could be found by repeating this step with other $\theta_{i}$.
To construct the training set for PSNN, different values of $\theta_{i}^{\prime}$ were sampled from each $\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ and solved by expanding the coefficient matrix, $\textbf{J}_{\mathcal{B_{i}}}$ with the corresponding binding constraints, running the calculations in eq. 13 and checking if the solutions satisfy the KKT conditions. This step was repeated using one other $\theta_{j}$ and fixing all other parameters, assuming the same $\textbf{CR}_{i}^{-},\textbf{CR}_{i}^{+}]$ would apply to this range. If the calculated solution does not satisfy KKT, another number was sampled from the same interval.